Problem 4.23. Suppose you take a sequence a Bernoulli trials with probability p = 0.4 of success. Let X be the number of trials you need to make to get 4 successful trials. Find EX and Var X
Answer #1

Here x is the number of trials require to get 4 succeful trials.

so here the distribution is negative binomial. with pdf

f(x) = (x-1)C3 (0.4)4(0.6)x-4

Here

E[X] = (1-p) r/p + 4 = 0.6 * 4/0.4 + 4 = 10

Var[X] = (1-p)r/p2 = 0.6 * 4/ 0.42= 15